Enoshima Company's leisure surfboat division manufactures surfboat fins and sells to group of suffers. Kenji Tanaka, the manager of the division has the following projections for the year 2016:

Sales (1,000 surfboat fins) $18,000

Operating expanses:

Variable expenses $10,500

Fixed expenses 4,500

Total expenses (15,000)   

Net Income $3,000

1. What is breakeven point of quantity sale?

2. If sales decrease 30%, what is loss or profit for a company?

3. Show the above relationship in graph. Cleary indicate the breakeven point and the situation described in question #2.
Answer #1

1)Breakeven point in quantity= fixed costs/(sales -variable cost per unit)
=4500/((18000-10500)/1000)
=600 units
2)sales decrease by 30% means new sales units=1000-(30%*1000)=700
sales per unit=18000/1000=18
variable cost per unit=10500/1000=10.5
New profit=sales-variable cost-fixed expenses
=(700*(18-10.5))-4500
=750
3)